Solve each system by addition, if possible.
x + 2y = -9
x - 2y = -1
------------------

Found 2 solutions by Alan3354, kekestarme:
Answer by Alan3354(69443)


x + 2y = -9
x - 2y = -1
------------------ Add them

Answer by kekestarme(2)

Copy and paste the links below to get a full description and CORRECT answer for your question.



First we need to know that the equation: y=mx+b



Go To: http://www.algebra.com/tutors/describe_linear_equation.solver?solver_action=plug for a better understanding & for the answer.

**I realized that both links are the same, just punch in the proper numbers to solve**

Go To: http://www.algebra.com/tutors/describe_linear_equation.solver?solver_action=plug for a better understanding & for the answer for x-2y=-1. **Remember that x-2y=-1 also equals x+-2y=-1 so don't get that mixed up because they are the same thing but just written in a different way.




If you need more help with problem(s) email: candyhelp411@gamil.com to get help! Also, check out my site: doyourstyle.webs.com where you can basically... DO YOU! Email me for anything else. Happy homeworking! :)